# UMGAD: Unsupervised Multiplex Graph Anomaly Detection

### v2

@@ IV-E Unsupervised Anomaly Score Threshold Selection Strategy
For the unsupervised graph anomaly detection task, since the nodes’ label (normal or abnormal) is unknown, a common approach is to compute the anomaly score for each node and rank them in order of the anomaly score from high to low. Nodes with high anomaly scores are considered anomalous nodes. However, determining a more accurate anomaly score threshold has always been a difficult problem in UGAD.
@@
$\Delta_{1}(i)=\bar{s}(i)-\bar{s}(i+1),\quad i=1,2,\dots,|\mathcal{V}|-w,$ (21)
@@
which corresponds to the region where the anomaly score sequence transitions from a steep decline to a plateau. The anomaly score threshold $s(\mathcal{T})$ is then defined as the smoothed score at the inflection point $\mathcal{T}$. If there exists several selectable points consistent with Eq. 23, we select the one with the smallest difference from $\bar{s}(|\mathcal{V}|)$ as $\mathcal{T}$. Therefore, nodes with scores $s(i)\geq s(\mathcal{T})$ are considered anomalous, while those with scores $s(i)<s(\mathcal{T})$ are deemed normal.
@@ IV-F Time complexity Analysis
The proposed UMGAD consists of three critical components: original-view graph reconstruction, augmented-view graph reconstruction, and dual-view contrastive learning. In original-view graph reconstruction, the time complexity of GMAE is $O(|\mathcal{V}|\times L\times f)$, and the time complexity of attention aggregation operation is $O(|\mathcal{V}|\times d_{h}\times f\times R)$, where $|\mathcal{V}|$ denotes the number of nodes, and $L$ represents the number of Simplified GCN layers, $d_{h}$ and $f$ denote the dimension of latent vectors and the size of node attribute, $R$ is the number of relational subgraphs. The overall time complexity of original-view graph reconstruction is $O(|\mathcal{V}|\times f\times(L+d_{h}\times R)$. For augmented-view graph reconstruction, the overall time complexity is $O(|\mathcal{V}|\times f\times(L+d_{h}\times R)$. The time complexity for dual-view contrastive learning is $O(|\mathcal{V}|\times f^{2}+|\mathcal{V}|\times f)$. Hence, the entire time complexity of the proposed UMGAD is approximately $O(|\mathcal{V}|\times f\times(L+d_{h}\times R+f)$.
@@ V Experiments
  • (RQ3) What are the effects of different modules in UMGAD on performance?
  • (RQ4) How do different hyperparameter settings affect the performance?
  • (RQ5) How does UMGAD perform compared to other GAD methods in the scenario of threshold selection with ground truth leakage?
  • (RQ6) How does UMGAD perform in terms of time efficiency compared to SOTA methods?
@@ V-A1 Dataset
We have conducted experiments on two datasets injected with synthetic anomalies: Retail_Rocket [47] (Retail for short) and Alibaba [48], and two real-world publicly available datasets with anomalies: Amazon [5] and YelpChi [49]. The detailed description of these datasets is shown as follows. Three datasets are adopted for evaluation:
  • Retail_Rocket (Retail for short). This is a benchmark dataset collected from the Retail_rocket recommendation system. In this dataset, interactive relations between users and items include page views (View), add-to-cart (Cart), and transactions (Buy).
  • Alibaba. This dataset is collected from Alibaba, one of the largest e-commerce platforms in China. It contains three interactive relations between users and items, i.e., page view (View), add to cart (Cart), and purchase (Buy).
  • Amazon. This dataset is collected from Amazon, it contains three interactive relations among users: U-P-U links users reviewing at least one product, U-S-U links users having at least one identical star rating within a week, U-V-U links linguistically similar users.
  • YelpChi. This dataset is collected from Yelp, it contains three interactive relations among users: R-U-R connects the reviews posted by the same user, R-S-R connects reviews that have the same identical star rating, R-T-R connects reviews that posted in the same term.
@@
YelpChi 45,954 6674 (R) R-U-R R-S-R R-T-R 49,315 3,402,743 573,616
Specifically, for synthetic anomalies, we follow the approach from [8] to introduce both structural and attribute anomalies. When injecting structural anomalies, we specify that a clique of size $m$ is introduced, i.e., if $m$ nodes are randomly selected from the network and then these nodes are made fully connected, all $m$ nodes in the clique are considered as anomalies. We iteratively repeat this process until $n$ factions are generated and the total number of structural anomalies is $m\times n$. Correspondingly, when injecting attribute anomalies, we first randomly select additional $m\times n$ nodes as attribute perturbation candidates. For each selected node $i$, we randomly select another $k$ nodes from the data and select the node $j$ whose attribute among the $k$ nodes deviates most from node $i$ by maximizing the Euclidean distance ${||x_{i}-x_{j}||}^{2}$ Then, we change the attribute $x_{i}$ of node $i$ to $x_{j}$. Additionally, for datasets containing real anomalies, we directly use the publicly available datasets from [9].
@@ V-A2 Baselines
We mainly compare five categories of methods on the unsupervised GAD task including the traditional method Radar [27].
@@
  • Radar [27]: It is a learning framework for characterizing the residuals of attribute information in anomaly detection and its consistency with network information.
Specific GAD Methods:
  • ComGA [28]: It is a community-aware anomaly detection framework that enhances GNNs through community-based segmentation and feature learning.
  • RAND [36]: It integrates reinforced neighborhood selection and improves the message-passing mechanism for unsupervised graph anomaly detection.
  • TAM [29]: It learns node embeddings for anomaly detection by maximizing local affinity and removing non-homogeneous edges to reduce bias.
CL-based GAD Methods:
  • CoLA [41]: It proposes a GNN-based contrastive model that captures high-dimensional attributes and local structures using a novel sampling method.
  • ANEMONE [50]: It employs a GNN encoder with multi-scale contrastive learning to capture graph patterns and a statistical estimator to detect node anomalies based on consistency.
  • Sub-CR [42]: It is a self-supervised framework combining multi-view comparison and attribute reconstruction to detect anomalies on attribute networks.
  • ARISE [51]: It is a graph anomaly detection model that identifies anomalies by learning substructures in attribute networks.
  • SL-GAD [52]: It uses contextual views with generative attribute regression and multi-view comparative learning to detect attribute, structural, and mixed anomalies.
  • PREM [53]: It includes a preprocessing module and an isochronous bit matching module to eliminate message-passing during training, using a simple contrastive loss to reduce training time and memory usage.
  • GCCAD [30]: It proposes a pre-training strategy to address label scarcity by generating pseudo anomalies through graph corruption.
  • GRADATE [31]: It is a multi-view, multi-scale contrastive learning framework that captures features to detect complex structural anomalies.
  • VGOD [32]: It is a variance-based graph outlier detection method that combines variance modeling and attribute reconstruction for balanced detection.
GAE-based GAD Methods:
  • GCNAE [54]: It is an unsupervised model using a variational graph autoencoder (VGAE) with GCN encoding and latent variables to learn graph embeddings.
  • DOMINANT [44]: It learns node embeddings using GCN to model graph topology and attributes, then reconstructs the original data with a deep autoencoder.
  • AnomalyDAE [55]: It is a joint representation learning model for anomaly detection using dual self-encoders to capture interactions between structure and node attributes.
  • AdONE [56]: It is an unsupervised AE approach with adversarial learning to separately learn network structure and attribute embeddings, minimizing the impact of outliers.
  • GAD-NR [45]: It combines neighborhood reconstruction for GAD problem, aiming to reconstruct a node’s entire neighborhood, including its local structure and attributes.
  • ADA-GAD [33]: It is a two-stage anomaly denoising autoencoder framework that reduces anomalies, trains a graph autoencoder, and applies node anomaly distribution regularization to prevent overfitting.
  • GADAM [57]: It is a local inconsistency mining approach that resolves conflicts with GNN messaging and enhances anomaly detection through improved message passing.
@@
  • AnomMAN [46]: It is a GCN-based framework for anomaly detection on multi-view attribute networks, using attention to fuse information across views.
  • DualGAD [13]: It is a self-supervised model combining generative and contrastive modules with masked subgraph reconstruction for improved node-level anomaly detection.
V-A3 Implementation Details
We implement all baselines according to their provided code or utilizing the PyGOD toolkit. We set the epoch number, dropout rate, and weight decay to 20, 0.1, and 0.01, respectively, and the embedding dimension $d$ to 32, and our method uses Simplified GCN as the encoder and decoder. For Amazon and YelpChi with real anomalies, the number of encoder layers is set to 2 and the depth of the decoder is set to 1. For Retail and Alibaba with synthetic anomalies, the number of encoder and decoder layers are both set to 1. The pre-training epoch and the retention epoch are both set to 20. AUC (area under the ROC curve) and Macro-F1 are used as performance metrics. All experiments are performed on 4 RTX 4090 GPUs with 24GB memory.
(a) Amazon (#real anomalies: 821)
(b) YelpChi (#real anomalies: 6674)
(c) Retail (#synthetic anomalies: 300)
(d) Alibaba (#synthetic anomalies: 300)
TABLE II: Performance comparison of all models on four datasets in the real unsupervised scenario.
@@ V-A3 Implementation Details
Trad. Radar [IJCAI’17] 0.625$\pm$0.002 0.533$\pm$0.007 0.659$\pm$0.003 0.560$\pm$0.008 0.582$\pm$0.007 0.525$\pm$0.004 0.502$\pm$0.003 0.485$\pm$0.011
MPI ComGA [WSDM’22] 0.626$\pm$0.015 0.553$\pm$0.018 0.591$\pm$0.003 0.573$\pm$0.006 0.661$\pm$0.006 0.691$\pm$0.013 0.514$\pm$0.001 0.491$\pm$0.001
RAND [ICDM’23] 0.665$\pm$0.007 0.622$\pm$0.010 0.692$\pm$0.002 0.603$\pm$0.007 0.660$\pm$0.004 0.687$\pm$0.008 0.525$\pm$0.005 0.501$\pm$0.002
TAM [NeurIPS’24] 0.688$\pm$0.006 0.650$\pm$0.007 0.712$\pm$0.004 0.660$\pm$0.008 0.763$\pm$0.011 0.637$\pm$0.008 0.546$\pm$0.007 0.538$\pm$0.008
CL CoLA [TNNLS’21] 0.554$\pm$0.013 0.513$\pm$0.011 0.570$\pm$0.011 0.541$\pm$0.018 0.582$\pm$0.008 0.522$\pm$0.005 0.444$\pm$0.003 0.468$\pm$0.007
ANEMONE [CIKM’21] 0.630$\pm$0.004 0.589$\pm$0.009 0.651$\pm$0.002 0.581$\pm$0.008 0.622$\pm$0.013 0.550$\pm$0.014 0.503$\pm$0.011 0.492$\pm$0.018
Sub-CR [IJCAI’22] 0.650$\pm$0.006 0.594$\pm$0.008 0.652$\pm$0.017 0.599$\pm$0.012 0.599$\pm$0.005 0.530$\pm$0.007 0.489$\pm$0.012 0.490$\pm$0.013
ARISE [TNNLS’23] 0.671$\pm$0.011 0.635$\pm$0.017 0.708$\pm$0.019 0.619$\pm$0.011 0.692$\pm$0.009 0.612$\pm$0.006 0.532$\pm$0.002 0.503$\pm$0.004
SL-GAD [TKDE’21] 0.658$\pm$0.018 0.596$\pm$0.016 0.700$\pm$0.008 0.611$\pm$0.009 0.653$\pm$0.008 0.585$\pm$0.003 0.522$\pm$0.005 0.484$\pm$0.007
PREM [ICDM’23] 0.661$\pm$0.008 0.609$\pm$0.009 0.680$\pm$0.016 0.607$\pm$0.008 0.658$\pm$0.018 0.588$\pm$0.003 0.521$\pm$0.008 0.493$\pm$0.006
GCCAD [TKDE’22] 0.665$\pm$0.011 0.611$\pm$0.006 0.701$\pm$0.006 0.630$\pm$0.004 0.662$\pm$0.007 0.591$\pm$0.012 0.522$\pm$0.013 0.497$\pm$0.007
GRADATE [AAAI’23] 0.683$\pm$0.004 0.644$\pm$0.004 0.702$\pm$0.007 0.642$\pm$0.009 0.730$\pm$0.002 0.661$\pm$0.012 0.540$\pm$0.001 0.521$\pm$0.003
VGOD [ICDE’23] 0.680$\pm$0.002 0.639$\pm$0.003 0.710$\pm$0.013 0.644$\pm$0.015 0.724$\pm$0.001 0.702$\pm$0.006 0.538$\pm$0.009 0.510$\pm$0.002
GAE DOMINANT [arXiv’16] 0.621$\pm$0.009 0.577$\pm$0.011 0.609$\pm$0.008 0.549$\pm$0.012 0.621$\pm$0.005 0.581$\pm$0.007 0.500$\pm$0.005 0.490$\pm$0.005
GCNAE [SDM’19] 0.623$\pm$0.007 0.579$\pm$0.010 0.624$\pm$0.007 0.564$\pm$0.011 0.607$\pm$0.017 0.583$\pm$0.010 0.502$\pm$0.010 0.494$\pm$0.008
AnomalyDAE [ICASSP’20] 0.618$\pm$0.003 0.525$\pm$0.005 0.665$\pm$0.007 0.571$\pm$0.008 0.634$\pm$0.015 0.588$\pm$0.012 0.523$\pm$0.016 0.496$\pm$0.010
AdONE [WSDM’20] 0.621$\pm$0.012 0.565$\pm$0.010 0.641$\pm$0.004 0.579$\pm$0.004 0.649$\pm$0.008 0.601$\pm$0.009 0.520$\pm$0.005 0.505$\pm$0.003
GAD-NR [WSDM’24] 0.681$\pm$0.006 0.632$\pm$0.005 0.707$\pm$0.007 0.658$\pm$0.010 0.752$\pm$0.006 0.681$\pm$0.005 0.540$\pm$0.002 0.531$\pm$0.008
ADA-GAD [AAAI’24] 0.688$\pm$0.003 0.650$\pm$0.002 0.715$\pm$0.002 0.663$\pm$0.002 0.754$\pm$0.001 0.687$\pm$0.003 0.545$\pm$0.006 0.530$\pm$0.007
GADAM [ICLR’24] 0.690$\pm$0.008 0.644$\pm$0.008 0.712$\pm$0.001 0.656$\pm$0.002 0.756$\pm$0.004 0.687$\pm$0.008 0.545$\pm$0.005 0.535$\pm$0.009
MV AnomMAN [IS’23] 0.688$\pm$0.005 0.645$\pm$0.008 0.710$\pm$0.006 0.650$\pm$0.009 0.758$\pm$0.005 0.689$\pm$0.007 0.542$\pm$0.006 0.536$\pm$0.004
DualGAD [IS’24] 0.688$\pm$0.007 0.647$\pm$0.003 0.712$\pm$0.005 0.653$\pm$0.004 0.758$\pm$0.002 0.688$\pm$0.006 0.542$\pm$0.003 0.538$\pm$0.008
UMGAD 0.770$\pm$0.009 0.722$\pm$0.005 0.825$\pm$0.006 0.740$\pm$0.006 0.878$\pm$0.005 0.779$\pm$0.004 0.608$\pm$0.001 0.597$\pm$0.004

### v4

@@ IV-E Unsupervised Anomaly Score Threshold Selection Strategy
For the unsupervised graph anomaly detection task, since the nodes’ labels (normal or abnormal) are unknown, a common approach is to compute the anomaly score for each node and rank them in order of the anomaly score from high to low. Nodes with high anomaly scores are considered anomalous nodes. However, determining a more accurate anomaly score threshold has always been a difficult problem in UGAD.
@@
$\Delta_{1}(i)=\bar{s}(i)-\bar{s}(i+1),\quad i=1,2,\dots,|\mathcal{V}|-w.$ (21)
@@
which corresponds to the region where the anomaly score sequence transitions from a steep decline to a plateau. The anomaly score threshold $s(\mathcal{T})$ is then defined as the smoothed score at the inflection point $\mathcal{T}$. If there exist several selectable points consistent with Eq. (23), we select the one with the smallest difference from $\bar{s}(|\mathcal{V}|)$ as $\mathcal{T}$. Therefore, nodes with scores $s(i)\geq s(\mathcal{T})$ are considered anomalous, while those with scores $s(i)<s(\mathcal{T})$ are deemed normal.
@@ IV-F Time Complexity Analysis
The proposed UMGAD consists of three critical components: original-view graph reconstruction, augmented-view graph reconstruction, and dual-view contrastive learning. In original-view graph reconstruction, the time complexity of GMAE is $O(|\mathcal{V}|\times L\times f)$, and the time complexity of attention aggregation operation is $O(|\mathcal{V}|\times d_{h}\times f\times R)$, where $|\mathcal{V}|$ denotes the number of nodes, and $L$ represents the number of Simplified GCN layers, $d_{h}$ and $f$ denote the dimension of latent vectors during graph convolution and the size of node attribute, $R$ is the number of relational subgraphs. The overall time complexity of original-view graph reconstruction is $O(|\mathcal{V}|\times f\times(L+d_{h}\times R)$. For augmented-view graph reconstruction, the overall time complexity is $O(|\mathcal{V}|\times f\times(L+d_{h}\times R)$. The time complexity for dual-view contrastive learning is $O(|\mathcal{V}|\times f^{2}+|\mathcal{V}|\times f)$. Hence, the entire time complexity of the proposed UMGAD is approximately $O(|\mathcal{V}|\times f\times(L+d_{h}\times R+f)$.
@@ V Experiments
  • (RQ3) How does UMGAD perform on large-scale graphs?
  • (RQ4) What are the effects of different modules in UMGAD on performance?
  • (RQ5) How do different hyperparameter settings affect the performance?
  • (RQ6) How does UMGAD perform compared to other GAD methods in the scenario of threshold selection with ground truth leakage?
  • (RQ7) How does UMGAD perform in terms of time efficiency compared to SOTA methods?
@@ V-A1 Dataset
We have conducted experiments on two datasets with injected anomalies: Retail_Rocket [50] (Retail for short) and Alibaba [51], and four real-world publicly available datasets with anomalies: Amazon [52], YelpChi [53], DGraph-Fin (DG-Fin for short) [54] and T-Social [20]. The detailed description of these datasets is shown as follows.
  • Retail_Rocket (Retail for short) contains user-item interactions such as page views (View), add-to-cart (Cart), and transactions (Buy).
  • Alibaba includes three types of user-item interactions like Retail: View, Cart, and Buy.
  • Amazon contains three types of user interactions: U-P-U links users who have reviewed at least one common product, U-S-U links users with identical star ratings within a week, and U-V-U links linguistically similar users.
  • YelpChi contains three interactive relations among users: R-U-R connects the reviews posted by the same user, R-S-R connects reviews that have the same identical star rating, R-T-R connects reviews that posted in the same term.
  • DGraph-Fin (DG-Fin for short) contains three interactive relations among users: U-C-U connects the emergency contact among users, U-B-U connects the borrowing behavior among users, U-R-U indicates that users follow each other.
  • T-Social contains three interactive relations among users: U-R-U links users who are friends with each other, U-F-U links users having fraud or money laundering behavior, U-G-U links online gambling users.
@@
YelpChi 45,954 6,674 (R) R-U-R R-S-R R-T-R 49,315 3,402,743 573,616
DG-Fin 3,700,550 15,509 (R) U-C-U U-B-U U-R-U 441,128 2,474,949 1,384,922
T-Social 5,781,065 174,010 (R) U-R-U U-F-U U-G-U 67,732,284 3,025,679 2,347,545
For anomaly injection, we follow the method in [55] to introduce both structural and attribute anomalies. For structural anomalies, a clique of size $m$ is formed by randomly selecting $m$ nodes and fully connecting them with one or multiple randomly assigned relation types. All $m$ nodes in the clique are considered anomalies. This process is repeated iteratively until $n$ factions are created, yielding $m\times n$ structural anomalies. For attribute anomalies, we randomly select $m\times n$ nodes and for each, identify the most deviant attribute by maximizing the Euclidean distance ${||x_{i}-x_{j}||}^{2}$ between node $i$ and a randomly selected node $j$. We then update node $i$’s attribute to match node $j$’s. For datasets with real anomalies, we directly use publicly available datasets.
@@ V-A2 Baselines
We mainly compare five categories of methods on the unsupervised GAD task.
(a) Retail (#injected anomalies: 300)
(b) Alibaba (#injected anomalies: 300)
(c) Amazon (#real anomalies: 821)
(d) YelpChi (#real anomalies: 6,674)
(e) DG-Fin (#real anomalies: 15,509)
(f) T-Social (#real anomalies: 174,010)
@@
  • Radar [27] characterizes the residuals of attribute information and its consistency with network information.
Specific MPI Methods:
  • ComGA [28] enhances GNNs through community-based segmentation and feature learning.
  • RAND [39] integrates reinforced neighborhood selection and improves the message-passing mechanism for UGAD.
  • TAM [29] learns embeddings by maximizing local affinity and removing non-homogeneous edges to reduce bias.
CL-based Methods:
  • CoLA [44] proposes a GNN-based contrastive model that captures high-dimensional attributes and local structures.
  • ANEMONE [56] combines a multi-scale contrastive GNN encoder and a statistical consistency estimator.
  • Sub-CR [45] combines multi-view comparison and attribute reconstruction to detect anomalies on attribute networks.
  • ARISE [57] identifies anomalies by learning substructures in attribute networks.
  • SL-GAD [58] uses contextual views with generative attribute regression and multi-view comparative learning.
  • PREM [59] eliminates training-phase message passing through preprocessing and isochronous bit matching.
  • GCCAD [30] proposes a pre-training strategy that generates pseudo anomalies via graph corruption to face label scarcity.
  • GRADATE [31] is a multi-view, multi-scale contrastive learning model that detects complex structural anomalies.
  • VGOD [32] is a variance-based model that combines variance modeling and attribute reconstruction.
GAE-based Methods
  • GCNAE [60] is an unsupervised model using a VGAE with GCN encoding and latent variables to embed graphs.
  • DOMINANT [47] models graph topology and attributes using GCN and reconstructs the original graph with AE.
  • AnomalyDAE [61] uses dual self-encoders to capture interactions between structure and node attributes.
  • AdONE [62] is an unsupervised AE approach with adversarial learning to separately learn structure and attribute.
  • GAD-NR [48] combines neighborhood reconstruction to reconstruct a node’s entire neighborhood.
  • ADA-GAD [33] is a two-stage anomaly denoising autoencoder that reduces anomalies, trains a graph autoencoder, and applies node regularization to prevent overfitting.
  • GADAM [63] is a local inconsistency mining approach that resolves conflicts with GNN messaging.
@@
  • AnomMAN [49] focuses UGAD on multi-view attribute networks using attention to fuse information across views.
  • DualGAD [13] combines generative and contrastive modules with masked subgraph reconstruction for UGAD.
TABLE II: Performance comparison of all models on four small-scale datasets in the real unsupervised scenario.
@@
Trad. Radar [IJCAI’17] 0.625$\pm$0.002 0.533$\pm$0.007 0.659$\pm$0.003 0.560$\pm$0.008 0.582$\pm$0.007 0.485$\pm$0.004 0.502$\pm$0.003 0.383$\pm$0.011
MPI ComGA [WSDM’22] 0.626$\pm$0.015 0.553$\pm$0.018 0.591$\pm$0.003 0.473$\pm$0.006 0.661$\pm$0.006 0.558$\pm$0.013 0.535$\pm$0.001 0.391$\pm$0.001
RAND [ICDM’23] 0.665$\pm$0.007 0.622$\pm$0.010 0.692$\pm$0.002 0.603$\pm$0.007 0.660$\pm$0.004 0.550$\pm$0.008 0.553$\pm$0.005 0.401$\pm$0.002
TAM [NeurIPS’24] 0.676$\pm$0.006 0.630$\pm$0.007 0.712$\pm$0.004 0.650$\pm$0.008 0.763$\pm$0.011 0.643$\pm$0.008 0.570$\pm$0.007 0.405$\pm$0.008
CL CoLA [TNNLS’21] 0.554$\pm$0.013 0.513$\pm$0.011 0.570$\pm$0.011 0.441$\pm$0.018 0.582$\pm$0.008 0.488$\pm$0.005 0.764$\pm$0.003 0.503$\pm$0.007
ANEMONE [CIKM’21] 0.630$\pm$0.004 0.589$\pm$0.009 0.651$\pm$0.002 0.581$\pm$0.008 0.622$\pm$0.013 0.510$\pm$0.014 0.773$\pm$0.011 0.502$\pm$0.018
Sub-CR [IJCAI’22] 0.650$\pm$0.006 0.594$\pm$0.008 0.652$\pm$0.017 0.599$\pm$0.012 0.599$\pm$0.005 0.493$\pm$0.007 0.781$\pm$0.012 0.513$\pm$0.013
ARISE [TNNLS’23] 0.671$\pm$0.011 0.623$\pm$0.017 0.708$\pm$0.019 0.619$\pm$0.011 0.692$\pm$0.009 0.579$\pm$0.006 0.779$\pm$0.002 0.509$\pm$0.004
SL-GAD [TKDE’21] 0.658$\pm$0.018 0.596$\pm$0.016 0.700$\pm$0.008 0.611$\pm$0.009 0.653$\pm$0.008 0.545$\pm$0.003 0.798$\pm$0.005 0.521$\pm$0.007
PREM [ICDM’23] 0.661$\pm$0.008 0.609$\pm$0.009 0.680$\pm$0.016 0.607$\pm$0.008 0.658$\pm$0.018 0.547$\pm$0.003 0.765$\pm$0.008 0.503$\pm$0.006
GCCAD [TKDE’22] 0.665$\pm$0.011 0.611$\pm$0.006 0.701$\pm$0.006 0.630$\pm$0.004 0.662$\pm$0.007 0.550$\pm$0.012 0.772$\pm$0.013 0.507$\pm$0.007
GRADATE [AAAI’23] 0.683$\pm$0.004 0.644$\pm$0.004 0.702$\pm$0.007 0.642$\pm$0.009 0.730$\pm$0.002 0.616$\pm$0.012 0.750$\pm$0.001 0.501$\pm$0.003
VGOD [ICDE’23] 0.680$\pm$0.002 0.639$\pm$0.003 0.710$\pm$0.013 0.644$\pm$0.015 0.724$\pm$0.001 0.606$\pm$0.006 0.748$\pm$0.009 0.502$\pm$0.002
GAE DOMINANT [arXiv’16] 0.621$\pm$0.009 0.577$\pm$0.011 0.609$\pm$0.008 0.549$\pm$0.012 0.621$\pm$0.005 0.531$\pm$0.007 0.519$\pm$0.005 0.380$\pm$0.005
GCNAE [SDM’19] 0.623$\pm$0.007 0.579$\pm$0.010 0.624$\pm$0.007 0.564$\pm$0.011 0.617$\pm$0.017 0.534$\pm$0.010 0.520$\pm$0.010 0.385$\pm$0.008
AnomalyDAE [ICASSP’20] 0.618$\pm$0.003 0.545$\pm$0.005 0.665$\pm$0.007 0.601$\pm$0.008 0.634$\pm$0.015 0.538$\pm$0.012 0.573$\pm$0.016 0.400$\pm$0.010
AdONE [WSDM’20] 0.621$\pm$0.012 0.565$\pm$0.010 0.641$\pm$0.004 0.579$\pm$0.004 0.649$\pm$0.008 0.561$\pm$0.009 0.670$\pm$0.005 0.497$\pm$0.003
GAD-NR [WSDM’24] 0.681$\pm$0.006 0.632$\pm$0.005 0.707$\pm$0.007 0.638$\pm$0.010 0.754$\pm$0.006 0.636$\pm$0.005 0.790$\pm$0.002 0.519$\pm$0.008
ADA-GAD [AAAI’24] 0.682$\pm$0.003 0.645$\pm$0.002 0.715$\pm$0.002 0.663$\pm$0.002 0.754$\pm$0.001 0.639$\pm$0.003 0.799$\pm$0.006 0.535$\pm$0.007
GADAM [ICLR’24] 0.690$\pm$0.008 0.644$\pm$0.008 0.705$\pm$0.001 0.636$\pm$0.002 0.752$\pm$0.004 0.633$\pm$0.008 0.796$\pm$0.005 0.533$\pm$0.009
MV AnomMAN [IS’23] 0.688$\pm$0.005 0.647$\pm$0.008 0.710$\pm$0.006 0.650$\pm$0.009 0.755$\pm$0.005 0.645$\pm$0.007 0.801$\pm$0.006 0.530$\pm$0.004
DualGAD [IS’24] 0.689$\pm$0.007 0.650$\pm$0.003 0.712$\pm$0.005 0.653$\pm$0.004 0.758$\pm$0.002 0.643$\pm$0.006 0.814$\pm$0.003 0.545$\pm$0.008
UMGAD 0.770$\pm$0.009 0.722$\pm$0.005 0.825$\pm$0.006 0.740$\pm$0.006 0.878$\pm$0.005 0.729$\pm$0.004 0.908$\pm$0.001 0.605$\pm$0.004
@@
V-A3 Implementation Details
We implement all baselines according to their provided code or by utilizing the PyGOD toolkit. Our method adopts GAT and simplified GCN as the encoder and decoder. For Amazon and YelpChi with real anomalies, the depth of the encoder and decoder is set to 2 and 1. For Retail and Alibaba with injected anomalies, the number of encoder and decoder layers is set to 1. AUC (area under the ROC curve) and Macro-F1 are used as performance metrics. All experiments are performed on the RTX 4090 GPU with 24GB of memory.